REPOST: Height of a post. Betty observed that the lamp post in front of her house casts a shadow of length 8 feet when the angle of inclination of the sun is 60 degrees. How tall is the lamp post? (In a 30-60-90 right triangle, the side opposite 30 is one-half the length of the hypotenuse.)
Answer by saltzy57(8)

Remember Pythagorean Theorem
a^2 + b^2 = c^2
side a = 8
side b = 1/2 side c
side c = hypotenuese
8^2 + 1/2*c^2 = c^2
64=1/2 c^2
128= c^2
c=sqrt (128)
b= 1/2 sqrt (128)
good luck